The displacement of the particle at x = 0 of a stretched string carrying a wave in the positive x-direction is given by f(t) = A sin(t/T). The wave speed is v. Write the wave equation.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. . At x = 0, f(t) = a sin (t/T) Wave speed = v ⇒ λ = wavelength = vT (T = Time period) So, general equation of wave Y = A sin [(t/T) – (x/vT)] [because y = f((t/T) – (x/ λ))